Determine the forces in members CG and GH of the symmetrically loaded truss. The forces are positive if in tension and negative if in compression.Question 19 View Policies Current Attempt in Progress Determine the forces in members CG and GH of the symmetrically loaded t
Answer #1

Given the symmetys Au= Fy=2+4=34A Now, cut a section left of a-a, and draw FED xam & is fod. Oftam (1/4)=26.560 BK en fa {fy=